# Orchestrator-Agent Trust: A Modular Agentic AI Visual Classification System with Trust-Aware Orchestration and RAG-Based Reasoning

### v2

@@ 2.3 Trust-Aware Orchestration with RAG and Re-Evaluation Loops
Confidence-Accuracy Calibration and Overconfidence Mitigation.
@@
3 Discussion
The increasing complexity and autonomy of AI systems calls for robust, interpretable, and generalizable architectures that can reason, evaluate, and adapt in real time [28]. This study presents a systematic exploration of a novel trust-aware agentic AI framework that blends zero-shot vision-language agents with orchestration, trust calibration, and RAG-based re-evaluation. Our three-tier experimental structure reveals key insights into how such systems can be structured for both performance and scalability while preserving explainability and adaptability.
A central insight from this study is the critical role of calibration and trust estimation in agentic AI systems. While traditional ensemble methods often aggregate agent outputs under the assumption of independent and reliable performance [29, 30], such strategies falter in real-world settings especially under zero-shot conditions where agents may exhibit systematic overconfidence or miscalibration [31, 32]. In these contexts, naïvely trusting self-reported confidence scores can lead to compounding errors, particularly in scenarios that demand high reliability and interpretability.
@@ 3 Discussion
Traditional ensemble learning methods, such as majority voting [33, 34], mean averaging or confidence-weighted fusion [35], operate under the assumption that model outputs are statistically independent and equally reliable. While effective in low-noise environments or when models are homogeneously calibrated, these techniques struggle in heterogeneous, zero-shot, or high-uncertainty scenarios where agent predictions may be misaligned or systematically overconfident [36, 37]. In contrast, our trust-aware orchestrator does not merely aggregate predictions it actively evaluates each agent’s trustworthiness using multi-dimensional metrics and selectively down-weights or discards predictions that exhibit poor calibration, inconsistency, or unjustified confidence. This approach aligns more closely with emerging frameworks in agentic AI, where orchestration involves structured reasoning across multiple autonomous agents with varying competencies.
@@
Prior work on dynamic task decomposition [38, 39], agent delegation [40, 41], and modular reasoning [42, 43] has shown that intelligent coordination across agents can outperform flat ensembles, particularly when agents contribute distinct skills or modalities [44]. However, many of these systems rely on rule-based or deterministic coordination logic and lack mechanisms for trust-based arbitration or reflective re-evaluation [45, 46]. Our orchestrator extends this space by integrating retrieval-augmented prompts and dynamic trust profiles, enabling recursive decision correction based on evidence-grounded feedback. Compared to static ensembles, this architecture enables real-time reasoning under uncertainty, improves robustness to adversarial disagreement, and supports scalable integration of new agents without retraining. Such orchestrator logic is increasingly critical as agent ecosystems grow in complexity and move toward plug-and-play, open-world operation.
@@
An additional strength of our framework lies in the re-evaluation loop powered by RAG. This feedback mechanism enables agents to reflect on their prior decisions in light of retrieved evidence from a vision-language knowledge base. By structuring the retrieved information around interpretable class definitions and visual prototypes, the system compensates for hallucination or semantic ambiguity common in zero-shot models. Our ablation analysis shows that although this loop was triggered in 100% of instances due to low trust profiling scores, disagreements between agents and system components reveal that GPT and Qwen often ignored re-evaluation prompts 12.5% and 29.38% of cases, respectively but this rarely led to correct predictions (only 3 out of 20 for GPT and 5 out of 47 for Qwen). These low success rates highlight the importance of the orchestrator, re-evaluation loop, and Image-RAG as authoritative sources, with optimal outcomes achieved through meta-reasoning arbitration. The practical implication is that retrieval-based grounding acts as an auxiliary supervision signal, enabling improvement without manual annotation or gradient updates.
@@
While RAG has shown considerable promise in enhancing large language models through external knowledge grounding especially in tasks such as medical decision support [47, 25], document retrieval and language alignment these systems typically lack a structured trust arbitration mechanism [26]. Most RAG implementations retrieve top-k textual or visual exemplars to refine responses, but treat the retrieval step as static and apply equal weight to all retrieved content [48, 49, 50]. In contrast, our Image-RAG pipeline integrates visual retrieval with dynamic trust scoring, enabling iterative re-evaluation loops where low-confidence or conflicting predictions trigger a targeted grounding process. This allows the system to not only retrieve relevant cases but also modulate decision-making based on model reliability and retrieval quality. Existing document-grounded agents focus on improving factuality but do not incorporate agent-level trust profiling [51]. Moreover, few approaches fuse visual similarity with structured re-prompting guided by trust thresholds [52]. Our design enhances interpretability by showing not only what evidence was retrieved but why it was considered trustworthy. This feedback loop is essential for deploying agentic AI in settings where error introspection and evidence traceability matter, and it represents a step beyond static RAG toward trust-calibrated retrieval orchestration.
@@
As agentic AI systems scale to incorporate dozens or even hundreds of specialized agents each with different modalities, competencies, or domain knowledge per-agent fine-tuning becomes computationally impractical and operationally rigid [53, 54]. Our trust-aware orchestration framework addresses this by absorbing agent heterogeneity through dynamic reliability profiling and selective arbitration, thereby supporting generalization across agents without task-specific adaptation. This design also mirrors recent advancements in user interface navigation agents and vision-language action systems, where orchestration is driven by flexible, intent-driven coordination rather than static aggregation. By separating reasoning logic from perception modules and incorporating natural language justifications, our system remains interpretable and auditable, supporting both technical transparency and human oversight two pillars necessary for scalable and trustworthy multi-agent AI ecosystems.
@@ 3.1 Summary, Limitations, and Future Perspective
Scaling the framework to accommodate larger and more diverse populations of agents poses challenges related to agent management, specialization, and dynamic trust assessment. Developing methods to efficiently integrate and update agent trust profiles will be critical for maintaining system robustness.
@@
4 Methods
The system comprises three experimental setups.
Experiment I
Each agent receives a prompt containing a single input image (from the test set) and is tasked with predicting the most appropriate plant disease class. In addition to the predicted class label, agents are required to return (i) a confidence score in the range [0.0–1.0] and (ii) a natural language explanation justifying their decision (Appendix, Fig. A1). These three outputs classification, confidence, and rationale are compiled and forwarded to a non-visual orchestrator model (o3-mini-2025-01-31 [55]), which serves as a comparative reasoner. The orchestrator does not process images directly; instead, it evaluates the agents’ predictions and justifications in light of their associated confidence scores and produces a final classification decision through structured, confidence-aware reasoning (Appendix, Fig. A2).
Experiment II
  1. 1.
    1. (a)
    2. (b)
  2. 2.
    Qwen-2.5-VL-7B Fine-Tuning: To adapt the Qwen-2.5 Vision-Language 7B (VL-7B) model for our task, we fine-tuned the unsloth/Qwen2.5-VL-7B-Instruct-bnb-4bit model [56], a 4-bit quantized variant of the original pretrained Qwen2.5 model. This quantized version enables efficient loading and inference with reduced memory requirements, making it suitable for commodity GPU hardware. We employed Low-Rank Adaptation (LoRA) in 16-bit precision to inject trainable adapters into the model, thereby allowing effective fine-tuning without full dequantization of the backbone weights.
Experiment III
Trust Evaluation Layer. The Trust Evaluation Layer conducts offline trust profiling of the agents by quantifying their calibration, discriminative reliability, and consistency over a labeled dataset. Specifically, both agents were evaluated in a zero-shot inference setting on the training image set originally employed in Experiment II for fine-tuning, but here presented without ground-truth labels during evaluation. For each image, we recorded the predicted class, associated confidence score, and explanatory output provided by the agent. Using this comprehensive log of model behavior, we computed a suite of quantitative trustworthiness metrics to assess each agent’s predictive confidence alignment and overall reliability under zero-shot conditions.
  • Overconfidence Ratio (OCR): The OCR measures the proportion of incorrect predictions made with high confidence (e.g., confidence ¿ 0.9), capturing the extent to which models exhibit unjustified certainty. A high OCR value signals a critical misalignment between model confidence and actual correctness, often leading to erroneous decisions in confidence-based orchestration pipelines. In our trust-aware framework, OCR serves as a penalizing factor in agent arbitration, allowing the orchestrator to detect and downweight predictions from agents that frequently make confidently wrong decisions.
    measuring how often models are confidently wrong.
  • Consistency Gap (CG): The CG quantifies the divergence in a model’s predictions when exposed to semantically equivalent but syntactically varied prompts, serving as a proxy for reasoning stability under linguistic perturbations. A large CG indicates that the agent’s output is highly sensitive to prompt phrasing, undermining reliability and reproducibility in decision-making. Within our framework, CG is used as a trust signal to evaluate the robustness of agent responses, enabling the orchestrator to identify models prone to inconsistency and initiate re-evaluation when necessary.
    $\text{CG}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}[P_{i}^{(1)}\neq P_{i}^{(2)}],$
    assessing variation in outputs across prompt formulations.
$\mathbf{e}_{i}=\text{L2\_normalize}(\text{ViT}_{\text{CLIP}}(\mathbf{I}_{i})),\quad\text{sim}(\mathbf{e}_{i},\mathbf{e}_{j})=\cos(\theta_{ij})=\mathbf{e}_{i}\cdot\mathbf{e}_{j}$ (1)
[…]
$R_{k}=\arg\max_{k}\{\cos(\mathbf{e}_{q},\mathbf{e}_{i})\mid i\in[1,N]\}$ (2)
$\text{conf}(c)=\frac{\sum_{i\in R_{k},y_{i}=c}s_{i}}{\sum_{j\in R_{k}}s_{j}}$ (3)
  {"category": "scab", "confidence": 0.5005},
  {"category": "healthy", "confidence": 0.3996},
  {"category": "rust", "confidence": 0.0999}
@@ 4 Methods
Funding: Partial financial support for open access publication was provided by the Hellenic Academic Libraries Link (HEAL-Link). Additionally, this work was supported in part by the National Science Foundation (NSF) and the United States Department of Agriculture (USDA), National Institute of Food and Agriculture (NIFA), through the “Artificial Intelligence (AI) Institute for Agriculture” program under Award Numbers AWD003473 and AWD004595, and USDA-NIFA Accession Number 1029004 for the project titled “Robotic Blossom Thinning with Soft Manipulators.” Additional support was provided through USDA-NIFA Grant Number 2024-67022-41788, Accession Number 1031712, under the project “ExPanding UCF AI Research To Novel Agricultural EngineeRing Applications (PARTNER).”
Conflict of interest: The authors declare that they have no conflict of interest.
Ethics approval and consent to participate: Not applicable.
Consent for publication: Not applicable.
Data availability: The datasets used and/or analyzed during the current study are available from the corresponding author upon reasonable request or through the GitHub repository (Source Link).
Materials availability: Not applicable.
Code availability: All source code and software modules developed in this study are publicly available at: https://github.com/Applied-AI-Research-Lab/Orchestrator-Agent-Trust
Author contribution: Konstantinos I. Roumeliotis and Ranjan Sapkota contributed equally to this work as lead authors. Ranjan Sapkota served as the corresponding author. All authors were involved in the study conception, system design, experimental execution, and manuscript preparation. All authors reviewed, revised, and approved the final version of the manuscript.
If any of the sections are not relevant to your manuscript, please include the heading and write ‘Not applicable’ for that section.
Appendix: Supplementary Figures
Figure A1: AI agent prompt and response for disease classification. The prompt instructs a vision-language model to classify apple leaf diseases. The response includes the predicted category, natural language justification, self-reported confidence score, processing latency, and estimated computational cost, returned in a structured JSON format.
Figure A2: Orchestrator prompt and agentic response for decision arbitration. In the agentic AI system, the orchestrator receives classification outputs from multiple vision-language agents and synthesizes them into a single, trusted decision. The JSON-formatted response includes the selected class, rationale for agreement or disagreement, confidence score, processing time, and token-based inference cost.
Figure A3: Experiment III Trust-aware orchestration with RAG and iterative re-evaluation. This flow diagram illustrates the dynamic reasoning pipeline in which the orchestrator triggers CLIP-based retrieval (Image-RAG) and a re-evaluation loop. Updated agent responses are scored by trust metrics to produce a final decision, enabling improved accuracy and interpretability for plant disease classification.
Appendix: Supplementary Tables
Table A1: Experiment I: Confidence score distribution for zero-shot agents and the orchestrator. Descriptive statistics for confidence outputs from Qwen-2.5-VL, GPT-4o, and the orchestrator under the zero-shot configuration.
@@
Table A2: Experiment II: Comparison of fine-tuning settings for GPT-4o. Performance and cost summary for ResNet-50–optimized versus default hyperparameter configurations.
@@
GPT-4o (ResNet-50 tuned) 10 16 0.0088 1,778 47.53
GPT-4o (Default settings) 3 1 0.0617 1,652 13.09
Table A3: Experiment II: Confidence statistics under few-shot setting. Descriptive statistics of model confidence scores post fine-tuning for both agents and the orchestrator.
@@
Table A4: Experiment III: Trust metrics for zero-shot agents under trust-aware orchestration. Reliability, calibration, and overconfidence profiling for Qwen and GPT-4o prior to re-evaluation.
[…]
@@
[…]
Table A5: Experiment III: Confidence scores after trust-based re-evaluation. Post-reflection confidence statistics from all agents under RAG-enhanced orchestration.

### v3

@@ 3.4.1 Time and Calibration Performance Across Configurations
3.4.2 Confidence-Accuracy Calibration and Overconfidence Mitigation
@@
4 Discussion and Future Directions
The increasing complexity and autonomy of AI systems calls for robust, interpretable, and generalizable architectures that can reason, evaluate, and adapt in real time Longo et al. (2024). This study presents a systematic exploration of a novel trust-aware agentic AI framework that blends zero-shot vision-language agents with orchestration, trust calibration, and RAG-based re-evaluation. Our three-tier experimental structure reveals key insights into how such systems can be structured for both performance and scalability while preserving explainability and adaptability.
A central insight from this study is the critical role of calibration and trust estimation in agentic AI systems. While traditional ensemble methods often aggregate agent outputs under the assumption of independent and reliable performance MacKenzie and Munster (2019); Ganaie et al. (2022), such strategies falter in real-world settings especially under zero-shot conditions where agents may exhibit systematic overconfidence or miscalibration Frei and Isotta (2019); Ojha et al. (2025). In these contexts, naïvely trusting self-reported confidence scores can lead to compounding errors, particularly in scenarios that demand high reliability and interpretability.
@@ 4 Discussion and Future Directions
Traditional ensemble learning methods, such as majority voting Singh et al. (2019); Yang et al. (2023), mean averaging or confidence-weighted fusion Brown (2017), operate under the assumption that model outputs are statistically independent and equally reliable. While effective in low-noise environments or when models are homogeneously calibrated, these techniques struggle in heterogeneous, zero-shot, or high-uncertainty scenarios where agent predictions may be misaligned or systematically overconfident Jahan et al. (2025); He and Jiang (2023). In contrast, our trust-aware orchestrator does not merely aggregate predictions it actively evaluates each agent’s trustworthiness using multi-dimensional metrics and selectively down-weights or discards predictions that exhibit poor calibration, inconsistency, or unjustified confidence. This approach aligns more closely with emerging frameworks in agentic AI, where orchestration involves structured reasoning across multiple autonomous agents with varying competencies.
@@
Prior work on dynamic task decomposition Flores Romero et al. (2025); Gao et al. (2024), agent delegation Fernández Domingos et al. (2022); Pataranutaporn et al. (2021), and modular reasoning Lu et al. (2025); Odobesku et al. (2025) has shown that intelligent coordination across agents can outperform flat ensembles, particularly when agents contribute distinct skills or modalities Liu et al. (2025). However, many of these systems rely on rule-based or deterministic coordination logic and lack mechanisms for trust-based arbitration or reflective re-evaluation Kermansaravi et al. (2025); Li et al. (2025). Our orchestrator extends this space by integrating retrieval-augmented prompts and dynamic trust profiles, enabling recursive decision correction based on evidence-grounded feedback. Compared to static ensembles, this architecture enables real-time reasoning under uncertainty, improves robustness to adversarial disagreement, and supports scalable integration of new agents without retraining. Such orchestrator logic is increasingly critical as agent ecosystems grow in complexity and move toward plug-and-play, open-world operation.
@@
An additional strength of our framework lies in the re-evaluation loop powered by RAG. This feedback mechanism enables AI agents to reflect on their prior decisions in light of retrieved evidence from a vision-language knowledge base. By structuring the retrieved information around interpretable class definitions and visual prototypes, the system compensates for hallucination or semantic ambiguity common in zero-shot models. Our ablation analysis shows that although this loop was triggered in 100% of instances due to low trust profiling scores, disagreements between agents and system components reveal that GPT and Qwen often ignored re-evaluation prompts 12.5% and 29.38% of cases, respectively but this rarely led to correct predictions (only 3 out of 20 for GPT and 5 out of 47 for Qwen). These low success rates highlight the importance of the orchestrator, re-evaluation loop, and Image-RAG as authoritative sources, with optimal outcomes achieved through meta-reasoning arbitration. The practical implication is that retrieval-based grounding acts as an auxiliary supervision signal, enabling improvement without manual annotation or gradient updates.
@@
While RAG has shown considerable promise in enhancing large language models through external knowledge grounding especially in tasks such as medical decision support Wada et al. (2025); Ke et al. (2025), document retrieval and language alignment these systems typically lack a structured trust arbitration mechanism Dong et al. (2025). Most RAG implementations retrieve top-k textual or visual exemplars to refine responses, but treat the retrieval step as static and apply equal weight to all retrieved content Yang et al. (2025); Zhang et al. (2025); Prince et al. (2024). In contrast, our Image-RAG pipeline integrates visual retrieval with dynamic trust scoring, enabling iterative re-evaluation loops where low-confidence or conflicting predictions trigger a targeted grounding process. This allows the system to not only retrieve relevant cases but also modulate decision-making based on model reliability and retrieval quality. Existing document-grounded agents focus on improving factuality but do not incorporate agent-level trust profiling Hammane et al. (2024). Moreover, few approaches fuse visual similarity with structured re-prompting guided by trust thresholds Li et al. (2024). Our design enhances interpretability by showing not only what evidence was retrieved but why it was considered trustworthy. This feedback loop is essential for deploying agentic AI in settings where error introspection and evidence traceability matter, and it represents a step beyond static RAG toward trust-calibrated retrieval orchestration.
@@
As agentic AI systems scale to incorporate dozens or even hundreds of specialized agents each with different modalities, competencies, or domain knowledge per-agent fine-tuning becomes computationally impractical and operationally rigid Ma et al. (2024a); Zhou et al. (2024). Our trust-aware orchestration framework addresses this by absorbing agent heterogeneity through dynamic reliability profiling and selective arbitration, thereby supporting generalization across agents without task-specific adaptation. This design also mirrors recent advancements in user interface navigation agents and vision-language action systems, where orchestration is driven by flexible, intent-driven coordination rather than static aggregation. By separating reasoning logic from perception modules and incorporating natural language justifications, our system remains interpretable and auditable, supporting both technical transparency and human oversight two pillars necessary for scalable and trustworthy multi-agent AI ecosystems.
@@ 4.1.3 Future Perspective
Scaling the framework to accommodate larger and more diverse populations of AI agents poses challenges related to agent management, specialization, and dynamic trust assessment. Developing methods to efficiently integrate and update agent trust profiles will be critical for maintaining system robustness.
@@
Acknowledgement
The publication of the article in OA mode was financially supported by HEAL-Link. This work was additionally supported in part by the National Science Foundation (NSF) and the United States Department of Agriculture (USDA), National Institute of Food and Agriculture (NIFA), through the “Artificial Intelligence (AI) Institute for Agriculture” program under Award Numbers AWD003473 and AWD004595, and USDA-NIFA Accession Number 1029004 for the project titled “Robotic Blossom Thinning with Soft Manipulators.” Additional support was provided through USDA-NIFA Grant Number 2024-67022-41788, Accession Number 1031712, under the project “Expanding UCF AI Research To Novel Agricultural Engineering Applications (PARTNER).”
@@ Acknowledgement
The authors declare no conflicts of interest.
Statement on AI Writing Assistance
ChatGPT and Grammarly were utilized to enhance grammatical accuracy and refine sentence structure; all AI-generated revisions were thoroughly reviewed and edited for relevance.
Appendix A. Supplementary Figures and Tables
Figure A1: AI agent prompt and response for disease classification. The prompt instructs a vision-language model to classify apple leaf diseases. The response includes the predicted category, justification, self-reported confidence, latency, and estimated computational cost in JSON format.
Figure A2: Orchestrator prompt and agentic response for decision arbitration. The orchestrator receives outputs from multiple agents and synthesizes them into a single trusted decision. The JSON response includes the class, rationale, confidence, processing time, and cost.
Figure A3: Experiment III: Trust-aware orchestration with RAG and re-evaluation. The orchestrator triggers CLIP-based retrieval (Image-RAG) and a re-evaluation loop. Updated agent responses are scored by trust metrics to produce a final decision.
Table A1: Experiment I: Confidence score distribution for zero-shot agents and the orchestrator.
@@ Appendix A. Supplementary Figures and Tables
Table A2: Experiment II: Comparison of fine-tuning settings for GPT-4o.
@@
GPT-4o (ResNet-50 tuned) 10 16 0.0088 1778 47.53
GPT-4o (Default settings) 3 1 0.0617 1652 13.09
Table A3: Experiment II: Confidence statistics under few-shot setting.
@@
Table A4: Experiment III: Trust metrics for zero-shot agents under trust-aware orchestration.
Model Acc. Avg. Conf. $_{\text{corr}} Conf$$_{\text{incorr}}$ CG OCR HCW THC CCC $p$-val ECE CWA
@@
Note: Acc.=Accuracy; OCR=Overconfidence Ratio; HCW=High-confidence wrong; THC=Total high-confidence; CCC=Confidence-Correctness Correlation; ECE=Expected Calibration Error; CWA=Confidence-Weighted Accuracy.
Table A5: Experiment III: Confidence scores after trust-based re-evaluation.